OpenStudy (anonymous):

I am having the craziest thought about calculus: So we know that \(\dfrac{d}{dx}x^n = nx^{n-1}\) and \(\dfrac{d^2}{dx^2}x^n = n(n-1)x^{n-2}\) So... What about \(\dfrac{d^{1/2}}{dx^{1/2}}x^n\)? Is it even possible?

OpenStudy (dan815):

haha good question

OpenStudy (anonymous):

@satellite73 @iambatman

OpenStudy (anonymous):

I believe degree should be \(n-\frac{1}{2}\), but that's just wild guess... What about coefficient? Guess nobody will know...

OpenStudy (anonymous):

@e.mccormick @freckles Any idea on how to start such thing?

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Fractional_calculus

OpenStudy (anonymous):

Looks scary... I mean I saw \(_aD_t^\alpha f(t)=\frac{d^n}{dt^n} {}_aD_t^{-(n-\alpha)}f(t)=\frac{d^n}{dt^n} {}_aI_t^{n-\alpha} f(t)\) Guess I am kind of not ready :(

OpenStudy (anonymous):

but dang I don't know it is already considered. lol

OpenStudy (anonymous):

Ye lol, I was asking myself: how would you define half a derivative? It has to be something close to half a power that you can use power rules to extend from integer powers. They really discuss this kind of thinking in the Heuristics section =)